OpenStudy (anonymous):

I need help ASAP (Elimination method) Not looking for just answers! I would like someone to work it through with me if possible! Use ELIMINATION method to solve: (x/2+y/3=7/5),(2x/3+3y/4=5/4)

OpenStudy (anonymous):

Ok just a moment I'll try to help you

OpenStudy (aum):

It would be easier if you get rid of the fractions by multiplying the equations throughout by the LCM.

OpenStudy (anonymous):

Yep it would

OpenStudy (anonymous):

So 12 for both?

OpenStudy (tkhunny):

No. You'll have to go to 60 to get them all!

OpenStudy (anonymous):

@tkhunny ok ill do that hold on plz!

OpenStudy (aum):

30 times the first equation 12 times the second equation

OpenStudy (aum):

30 times the first equation: 15x + 10y = 42 12 times the second equation: 8x + 9y = 15 Pick which variable you want to eliminate.

OpenStudy (anonymous):

@aum do you mean like this?

OpenStudy (aum):

No you are multiply the top by 30 and the bottom also by 30. But in an equation we can multiply the equation THROUGHOUT by any number and the equation will still be valid. For example, the equation, 2x = 4 remains the same if we multiply throughout by 5: 10x = 20 is same as 2x = 4. In both cases x = 2. So multiply the first equation throughout just by 30 because the least common denominator of 2, 3 and 5 is 30.

OpenStudy (aum):

Multiply the second equation throughout by 12 because the least common denominator of 3, 4 and 4 is 12.

OpenStudy (aum):

The reason we are multiplying is to get rid of the denominator and so the equations become easier to work with.

OpenStudy (anonymous):

Sorry I am like really confused.. isn't that what I did?

OpenStudy (aum):

\(\Large 30 * \frac x2 = 15x \) What you are doing is: \(\Large \frac{30}{30} * \frac x2 = \frac{30x}{60} \)

OpenStudy (anonymous):

ohhh ok so I am only multiplying the top! okayyy, hold on plz lol

OpenStudy (anonymous):

so like this?

OpenStudy (aum):

You got the LHS of both equations correct. But RHS of both are incorrect.

OpenStudy (anonymous):

What is RHS?

OpenStudy (aum):

RHS - Right Hand Side LHS _ Left Hand Side (of equations)

OpenStudy (anonymous):

ohh ok cool. so the 8x+9y=5 is incorrect?

OpenStudy (aum):

15x + 10y (the LHS) is correct. = 35 (RHS) is wrong. 8x + 9y (the LHS_ is correct. = 5 (the RHS) is wrong.

OpenStudy (aum):

\(\large \frac 75 * 30 = 42\) (not 35) \(\large \frac 54 * 12 = 15\) (not 5)

OpenStudy (anonymous):

oh lol im so dumb! Ok now I see what i did wrong! So i am going to finish the problem and ill post what i got and would you mind telling me if I did it correctly??

OpenStudy (aum):

sure.

OpenStudy (anonymous):

So either I am just reallllllly bad at math or I did something wrong... or just both ... @aum this is what I got but my math book says the answer is (3,-1) and I didn't get that...

OpenStudy (aum):

You may have made a mistake in the numbers in the original problem posting. Can you post a screenshot of the question?

OpenStudy (anonymous):

yep

OpenStudy (aum):

y = -111/55 is CORRECT for this problem as it appears in the original posting. (BTW, the last line in your answer should be y = -2.01 and not x = -2.01 ).

OpenStudy (aum):

Will you be posting a screenshot of the question?

OpenStudy (anonymous):

Oh ok lol ! an it is attached to this

OpenStudy (aum):

The mistake in your original posting is, the RHS of the first equation should be 7/6 and not 7/5.

OpenStudy (anonymous):

Ohhhh omg Sorry lol!! well at least I know how to do it now! So instead of multiplying by 30 I would do it by 6?

OpenStudy (aum):

correct. I will do it below from start to finish.

OpenStudy (aum):

x/2 + y/3 = 7/6 (multiply by 6) 3x + 2y = 7 ----- (1) 2x/3 + 3y/4 = 5/4 (multiply by 12) 8x + 9y = 15 ----- (2) 8 x (1): 24x + 16y = 56 -3 x (2): -24x - 27y = -45 add -11y = 11 y = -1 put it in (1) 3x + 2(-1) = 7 3x - 2 = 7 3x = 9 x = 3 x = 3, y = -1 or (3,-1)

OpenStudy (anonymous):

Awesome!!! Thank you so much for taking the time to work the problem out with me!

OpenStudy (aum):

You are welcome. :)